# Efficient Test Data Generation for MC/DC with OCL and Search

### v1

@@ 4.3.2 Constraint with Multiple Clauses
$A_{min}\mathrm{=}\left\{\begin{array}[]{l}(c_{min}-c_{min})*s_{f}\mathrm{,\ \ }c\mathrm{>}0\\ (c_{min}+c_{min})*s_{f}\mathrm{,\ \ }c\mathrm{<}0\\ -1*s_{f}\mathrm{,\ \ }c\mathrm{=}0\end{array}\right.$ (6)
$A_{max}\mathrm{=}\left\{\begin{array}[]{l}(c_{max}+c_{max})*s_{f}\mathrm{,\ \ }c\mathrm{>}0\\ (c_{max}-c_{max})*s_{f}\mathrm{,\ \ }c\mathrm{<}0\\ 1*s_{f}\mathrm{,\ \ }c\mathrm{=}0\end{array}\right.$ (7)
For example, consider a constraint C7 as shown in Listing 1. There are two constants i.e., $c=\{100,5000\}$ for the attribute flightDistance in this constraint. The minimum and maximum values in the constants collection c are $c_{min}=100$ and $c_{max}=5000$ respectively. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values for the range are $A_{min}=100-100*1=0$ and $A_{max}=5000+5000*1=10000$. Therefore, the range of the attribute flightDistance is $[0,10000]$.
@@
For example, consider a constraint C8 as shown in Listing 1. The constraint C8 contains an invariant with two clauses i.e., ‘self.mission.flightDistance $>$ self.uav.MIN_RANGE+100 and self.mission.flightDistance $>$ self.uav.MAX_RANGE-100’. In this clause, flightDistance is the attribute (A) and there are three different constants i.e., $c_{1}=$MIN_RANGE, $c_{2}=100$, and $c_{3}=$MAX_RANGE. Both clauses contain constants with addition and subtraction arithmetic operators (i.e., + and -). If the value of the constant MIN_RANGE is 20 and MAX_RANGE is 1000, the resultant constants after solving both arithmetic expressions are $c_{1_{a}}=120$ and $c_{2_{a}}=900$. Therefore, the minimum and maximum values in the constants collection $c_{a}$ are $c_{min}=120$ and $c_{max}=900$. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, $A_{min}=120-120*1=0$ and $A_{max}=900+900*1=1800$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,1800]$.
@@
For example, consider a constraint C9 as shown in Listing 1. The constraint consists of a pre-condition with two clauses i.e., ‘self.distance $>$ self.MAX_RANGE and maxDist $>$ minDist’. Both clauses contain one attribute distance, two different variables (i.e., maxDist and minDist), and a constant MAX_RANGE. The range for the first clause is calculated by using Equations 4 and 5. If the value of MAX_RANGE is 10 and the value of the scaling factor is set to 1, the range of attribute (distance) with $s_{f}=1$ is $[0,10]$. In the case of the second clause i.e., ‘maxDist $>$ minDist’, the heuristic described in Case 2 of Section 4.3.1 is used to calculate the range for both variables. If the smallest positive integer value selected at random is 5 and the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the range of maxDist and minDist with $s_{f}=1$ is $[0,10]$.
Case 3: Multiple clauses joined with Boolean operators and involve attributes/variables on both sides of a relational operator. This case considers constraints with multiple clauses containing attributes/variables on both sides of a relational operator and all or some clauses are dependent [^3] A clause is dependent on another clause if the same attribute/variable is used in both clauses. . The dependent clauses contain some dependent attributes/variables and some dependent attributes/variables. For all independent attributes/variables, the heuristic is to choose the smallest positive integer value randomly in the range [0, 100]. For each dependent attribute/variable, the heuristic is to choose the smallest positive integer value randomly in the range [0, 100]. The selected Integer value for each dependent and independent attribute/variable is considered as a constant c. Each constant is used to determine the range of all attributes or variables using Equations 4 and 5.
@@ 5 Empirical Evaluation
In this section, we present the empirical evaluation of our proposed strategy with CBR and range reduction methods. The goal of empirical evaluation is to analyze the effectiveness of proposed methods (CBR and range reduction). For the evaluation, we selected the AVM that solves each constraint individually using the approach by Ali et al. [4]. We select AVM because the final MC/DC constraints contain several conjunctions, disjunctions, and negations for which AVM has shown better performance than (1+1) EA, as reported in previous studies [4, 6]. Moreover, experimental results in  [4, 6] show that local search algorithms (e.g., AVM) perform better as compared to global search algorithms (e.g., GA). Therefore, our experiment includes AVM as a search algorithm. We use Random Search (RS) as a comparison baseline. We evaluate each method of our strategy using AVM to have a fair comparison. The original AVM is termed as AVMo, CBR with AVM is termed as AVMc, and range reduction with AVM is termed as AVMr. The evaluation is also aimed at comparing the combination of CBR and range reduction with the individual methods and is thus termed AVMrc. Additionally, we compare our proposed methods (i.e., AVMc, AVMr, and AVMrc) for solving MC/DC constraints with relevant constraint solvers. The following are the research questions for our empirical evaluation.
@@
For RQ1, we primarily evaluate AVMc, AVMr, and AVMo in terms of success rates. Whenever a search algorithm solves a constraint, it is considered a success, and its success rate is increased. A 100% success rate of an algorithm corresponding to a constraint means that the algorithm solves the constraint in every run. In some cases, when two algorithms achieve the same success rate, comparing the algorithms based on only success rates may not be sufficient to analyze which algorithm is better. Therefore, we used iterations count as another measure. The number of iterations a search algorithm takes in solving a constraint is termed an iteration count. We also compare algorithms based on the time to solve OCL constraints. Regarding RQ2, we are interested in analyzing the effect of combining AVMc and AVMr (i.e., AVMrc) in solving MC/DC constraints. For this analysis, we also use success rates, iteration counts, and time as the evaluation metrics. In RQ3, we compare our proposed MC/DC constraints solving methods (AVMc, AVMr, and AVMrc) with relevant constraint solvers, considering the percentage of successfully solved constraints and the average time taken by different solvers.
@@ 5.1 Experiment Design and Settings
For the experiment, we used six cross-domain case studies with different complexity levels. Table 3 shows statistics for each case study. One of the case studies is industrial which represents the software system of a Ground Control Station (GCS) used to remotely control Unmanned Aerial Vehicles [86]. Two case studies, EU-Rental [35] and the Royal and Loyal [100, 17] are widely referred to in OCL literature and are considered as benchmark case studies [59]. The fourth case study represents the UML meta-model of State Machines [75]. The fifth case study is an artificial case study (Seat Belt System) that we developed inspired by various industrial case studies published in the literature [4, 6]. The final case study is industrial which represents a Student Registration System (SRS).
@@
Due to randomness in the search algorithms, we repeated each algorithm 100 times for all constraints. We used iterations (fitness evaluations) as a stopping criterion. We set the maximum number of iterations to 2000 (used earlier in [4]). The experiment is executed on 15 independent machines including five machines with quad-core 1.2 GHz processor and 1 GB RAM, five machines with core i5 1.7 GHz processor and 8 GB RAM, and five machines with core i7 3.2 GHz processor and 32 GB RAM. All algorithms for the same case study are executed on a group of machines with the same specifications.
We analyze experiment results according to the guidelines presented in [9]. Our primary goal is to analyze search algorithms based on success rates. This leads to the dichotomous type of data. For this purpose, we use Fisher’s exact test to analyze the statistical significance and calculate the Odds Ratios to check which algorithm is better. In some cases, two search algorithms may have the same success rate and the comparison results are not statistically significant. Therefore, we also used iterations count as a measure. In the case algorithms are compared based on iterations count, we use the Wilcoxon test for statistical significance and Vargha-Delaney Â${}_{12}$ measure [99] to analyze which one is better. We set the statistical significance level to 0.05 (i.e., $\alpha=0.05$) which is an accepted practice in the domain [4, 6].
For the comparison of our methods with existing constraint solvers (RQ3), we conducted an experiment using all six case studies, comprising a total of 129 OCL constraints. We selected two constraint solvers for this comparison: a CSP-based solver, UMLtoCSP [19], and a SMT+Search-based solver, PLEDGE [95]. We chose these two constraint solvers to ensure a fair comparison. Both constraint solvers take UML class diagrams and OCL constraints as input, without requiring the transformation of OCL constraints into an intermediate formalism like Alloy, which would introduce additional overhead for industrial constraints [4]. The constraint solver we used operates similarly, making it a suitable comparison. We executed this experiment on one machine with a 3.2 GHz processor, 32 GB RAM, and a Windows 10 operating system.
@@ 5.2 Discussion on Results
Fig. 3 shows the success rates achieved by AVMr, AVMc, AVMrc, AVMo, and RS for all case studies. Table 4 shows the results of the Fisher Exact test and the Wilcoxon test for statistical significance and the results of the effect size calculated using Odds Ratios and Vargha-Delaney’s Â${}_{12}$ measure. For the comparison based on time, Fig. 4 shows the time taken for each OCL constraint during 100 evaluations. Table 5 shows the results of the Wilcoxon test for statistical significance and the results of the effect size calculated using Vargha-Delaney’s Â${}_{12}$ measure. The columns in both tables represent a comparison between two algorithms based on statistical significance and effect size corresponding to each case study and its total constraints. The values presented in both tables (i.e., Table 4 and Table 5) refer to the number of times p-value$<$0.05 for the results of Fisher Exact and the Wilcoxon test. Similarly, for the effect size measures (Odds Ratios and Vargha-Delaney’s Â${}_{12}$), the tables show the number of times (including percentage fraction) an algorithm is better than the other one, i.e., Algorithm_1 outperforms ($>$) Algorithm_2. We show aggregated results due to a large number of comparisons among algorithms corresponding to all case studies. Table 6 presents results of comparison among AVMr, AVMc, AVMrc, UMLtoCPS, and PLEDGE based on the percentage of solved MC/DC constraints and the average time.
@@
To compare algorithms based on success rates, the box plots in Fig. 3 show that AVMr achieved a median success rate of $\approx$100for five case studies (i.e., GCS, EUR, RnL, SM, and SRS). Similarly, AVMc achieved $\approx$100median success rate for five case studies (i.e., EUR, RnL, SM, SBS, and SRS). AVMo achieved a median success rate of $\approx$100for three case studies (i.e., RnL, SM, and SRS). Whereas RS cannot achieve a 100% median success rate for any of the case studies.
For the statistical comparison between AVMr and AVMc, Table 4 shows that AVMc outperforms AVMr more than 50% of the time for all six case studies. In the case of comparison between AVMr and AVMo, AVMr is better than AVMo for $\approx$60of the time for GCS, EUR, RnL, and SBS case studies. While comparing AVMc and AVMo, AVMc is better than AVMo for more than $\approx$60of the constraints for GCS, EUR, RnL, and SBS case studies. Whereas for two case studies i.e., SM and SRS, AVMo is better than AVMc and AVMr for $\approx$60of the cases. If we compare AVMr, AVMc, and AVMo with RS, we can notice that all three algorithms outperform RS for more than $\approx$80of the cases for all six case studies.
Comparing AVMr and AVMc based on time, from Fig. 4 and Table 5, it is observed that both AVMr and AVMc take approximately an equal amount of time on average. For GCS, RnL, SM, and SRS case studies, AVMc is significantly better than AVMr. For the comparison between AVMr and AVMo, it is observed that AVMo is better for $\approx$60of the time. However, for EUR, SBS, and SRS case studies, AVMr is better than AVMo for $\approx$60of the constraints. In the case of the comparison between AVMc and AVMo, AVMc outperforms AVMo for GCS, SM, and SRS in $\approx$60of the constraints. For the three case studies i.e., EUR, RnL, and SBS, AVMo outperforms AVMc in $\approx$60of the cases. In the case of comparing AVMr, AVMc, and AVMo with RS, we can analyze that all three algorithms outperform RS for $\approx$50of the cases especially for EUR, RnL, SM, and SBS.
@@
From the box plots shown in Fig. 3, it is clear that AVMrc achieved a median success rate of $\approx$100for all six case studies. Whereas individually the AVMr and AVMc achieved $\approx$100median success rate for five case studies. Moreover, AVMo achieved a median success rate of $\approx$100for only three case studies.
@@
For the comparison based on time, the graph in Fig. 4 shows that AVMrc takes less time as compared to AVMr, AVMc, and AVMo. The statistical comparison results given in Table 5 show that AVMrc performs better than AVMo for more than $\approx$60of the constraints specifically for GCS, EUR, SBS, and SRS case studies. While comparing AVMrc and AVMc, it is noticed that AVMrc is better than AVMc for more than $\approx$70of the cases specifically for EUR, SBS, and SRS case studies. Whereas for the GCS, RnL, and SM case studies, AVMc is better than AVMrc for $\approx$60of the cases. In the comparison between AVMrc and AVMr, it is observed that AVMrc outperforms AVMr for more than $\approx$90of the time for all six case studies. Finally, for the comparison between AVMrc and RS, it is clear that AVMrc performs better than RS for $\approx$60of the cases specifically for EUR, RnL, SM, and SBS case studies.
@@
A thorough analysis of UMLtoCPS and PLEDGE results, we observed that the reason for unsolved MC/DC constraints is unsatisfiability. MC/DC constraints can have conflicting clauses that can lead these solvers to unsatisfiable scenarios. Given that these solvers are not specifically designed for solving MC/DC constraints and handling conflicting constraints, a substantial number of constraints remain unsolved. Moreover, PLEDGE operates in a way that it builds constraints’ syntax tree, and determines which clauses to solve using SMT and which ones with search. During the constraint-solving process, if the clauses to solve using SMT turn out to be conflicting, SMT marks the constraint as unsatisfiable, resulting in unsolved constraints.

### v2

@@ 4.3.2 Constraint with Multiple Clauses
$A_{min}\mathrm{=}\left\{\begin{array}[]{ l l }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}&\quad c\mathrm{>}0\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*c_{min}}*s_{f}&\quad c\mathrm{<}0\\ -1*s_{f}&\quad c\mathrm{=}0\end{array}\right.$ (6)
$A_{max}\mathrm{=}\left\{\begin{array}[]{ l l }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*c_{max}}*s_{f}&\quad c\mathrm{>}0\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}&\quad c\mathrm{<}0\\ 1*s_{f}&\quad c\mathrm{=}0\end{array}\right.$ (7)
For example, consider a constraint C7 as shown in Listing 1. There are two constant values, i.e., $c=\{100,5000\}$ for the attribute flightDistance in this constraint. The minimum and maximum values in the constants collection c are $c_{min}=100$ and $c_{max}=5000$ respectively. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values for the range are $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*5000*=10000}$. Therefore, the range of the attribute flightDistance is $[0,10000]$.
@@
For example, consider a constraint C8 as shown in Listing 1. The constraint C8 contains an invariant with two clauses, i.e., ‘self.mission.flightDistance $>$ self.uav.MIN_RANGE+100 and self.mission.flightDistance $>$ self.uav.MAX_RANGE-100’. In this clause, flightDistance is the attribute (A) and there are three different constants, i.e., $c_{1}=$MIN_RANGE, $c_{2}=100$, and $c_{3}=$MAX_RANGE. Both clauses contain constants with addition and subtraction arithmetic operators (i.e., + and -). If the value of the constant MIN_RANGE is 20 and MAX_RANGE is 1000, the resultant constants after solving both arithmetic expressions are $c_{1_{a}}=120$ and $c_{2_{a}}=900$. Therefore, the minimum and maximum values in the constants collection $c_{a}$ are $c_{min}=120$ and $c_{max}=900$. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*900*1=1800}$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,1800]$.
@@
For example, consider a constraint C9 as shown in Listing 1. The constraint consists of a pre-condition with two clauses, i.e., ‘self.distance $>$ self.MAX_RANGE and maxDist $>$ minDist’. Both clauses contain one attribute distance, two different variables (i.e., maxDist and minDist), and a constant MAX_RANGE. The range for the first clause is calculated by using Equations 4 and 5. If the value of MAX_RANGE is 10 and the value of the scaling factor is set to 1, the range of attribute (distance) with $s_{f}=1$ is $[0,10]$. In the case of the second clause, i.e., ‘maxDist $>$ minDist’, the heuristic described in Case 2 of Section 4.3.1 can be used to calculate the range for both variables. If the smallest positive integer value selected at random is 5 and the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the range of maxDist and minDist with $s_{f}=1$ is $[0,10]$.
Case 3: Multiple clauses joined with Boolean operators and involve attributes/variables on both sides of a relational operator. This case considers constraints with multiple clauses containing attributes/variables on both sides of a relational operator and all or some clauses are dependent [^3] A clause is dependent on another clause if the same attribute/variable is used in both clauses. . The dependent clauses contain some dependent attributes/variables and some dependent attributes/variables. For all independent attributes/variables, the heuristic is to choose the smallest positive integer value randomly in the range [0, 100]. For each dependent attribute/variable, the heuristic is to choose the smallest positive integer value randomly in the range [0, 100]. The selected Integer value for each dependent and independent attribute/variable is considered a constant c. Each constant determines the range of all attributes or variables using Equations 4 and 5.
@@ 5 Empirical Evaluation
In this section, we present the empirical evaluation of our proposed strategy with CBR and range reduction methods. The goal of empirical evaluation is to analyze the effectiveness of proposed methods (CBR and range reduction). For the evaluation, we selected the AVM that solves each constraint individually using the approach by Ali et al. [4]. We select AVM because the final MC/DC constraints contain several conjunctions, disjunctions, and negations for which AVM has shown better performance than (1+1) EA, as reported in previous studies by Ali et al. [4, 6]. Moreover, experimental results by Ali et al. [4, 6] show that local search algorithms (e.g., AVM) perform better as compared to global search algorithms (e.g., GA). Therefore, our experiment includes AVM as a search algorithm. We use Random Search (RS) as a comparison baseline. We evaluate each method of our strategy using AVM to have a fair comparison. The original AVM is termed AVMo, CBR with AVM is termed AVMc, and range reduction with AVM is termed AVMr. The evaluation is also aimed at comparing the combination of CBR and range reduction with the individual methods and is thus termed AVMrc. Additionally, we compare our proposed methods (i.e., AVMc, AVMr, and AVMrc) for solving MC/DC constraints with relevant constraint solvers. The following are the research questions for our empirical evaluation.
@@
For RQ1, we primarily evaluate AVMc, AVMr, and AVMo in terms of success rates. Whenever a search algorithm solves a constraint, it is considered a success, and its success rate is increased. A 100% success rate of an algorithm corresponding to a constraint means that the algorithm solves the constraint in every run. In some cases, when two algorithms achieve the same success rate, comparing the algorithms based on only success rates may not be sufficient to analyze which algorithm is better. Therefore, we used iterations count as another measure. The number of iterations a search algorithm takes in solving a constraint is termed an iteration count. We also compare algorithms based on the time to solve OCL constraints. Regarding RQ2, we are interested in analyzing the effect of combining AVMc and AVMr (i.e., AVMrc) in solving MC/DC constraints. For this analysis, we also use success rates, iteration counts, and time as the evaluation metrics. In RQ3, we compare our proposed MC/DC constraint-solving methods (AVMc, AVMr, and AVMrc) with relevant constraint solvers, considering the percentage of successfully solved constraints and the average time taken by different solvers.
@@ 5.1 Experiment Design and Settings
For the experiment, we used six cross-domain case studies with different complexity levels. Table 3 shows statistics for each case study. One of the case studies is industrial which represents the software system of a Ground Control Station (GCS) used to remotely control Unmanned Aerial Vehicles [86, 90]. Two case studies, EU-Rental [35] and the Royal and Loyal [100, 17] are widely referred to in OCL literature and are considered as benchmark case studies [59]. The fourth case study represents the UML meta-model of State Machines [75]. The fifth case study is an artificial case study (Seat Belt System) that we developed inspired by various industrial case studies published in the literature [4, 6]. The final case study is industrial which represents a Student Registration System (SRS).
@@
Due to randomness in the search algorithms, we repeated each algorithm 100 times for all constraints. We used iterations (fitness evaluations) as a stopping criterion. We set the maximum number of iterations to 2000 (used earlier by Ali et al. [4]). The experiment is executed on 15 independent machines including five machines with quad-core 1.2 GHz processor and 1 GB RAM, five machines with core i5 1.7 GHz processor and 8 GB RAM, and five machines with core i7 3.2 GHz processor and 32 GB RAM. All algorithms for the same case study are executed on a group of machines with the same specifications.
We analyze experiment results according to the guidelines presented by Arcuri et al. [9]. Our primary goal is to analyze search algorithms based on success rates. This leads to the dichotomous type of data. For this purpose, we use Fisher’s exact test to analyze the statistical significance and calculate the Odds Ratios to check which algorithm is better. In some cases, two search algorithms may have the same success rate and the comparison results are not statistically significant. Therefore, we also used iterations count as a measure. In the case algorithms are compared based on iterations count, we use the Wilcoxon test for statistical significance and Vargha-Delaney Â12 measure [99] to analyze which one is better. We set the statistical significance level to 0.05 (i.e., $\alpha=0.05$) which is an accepted practice in the domain [4, 6].
For the comparison of our methods with existing constraint solvers (RQ3), we conducted an experiment using all six case studies, comprising a total of 129 OCL constraints. We selected two constraint solvers for this comparison: a CSP-based solver, UMLtoCSP [19], and a SMT+Search-based solver, PLEDGE [95]. We chose these two constraint solvers to ensure a fair comparison. Both constraint solvers take UML class diagrams and OCL constraints as input, without requiring the transformation of OCL constraints into an intermediate formalism like Alloy, which would introduce additional overhead for industrial constraints [4]. The constraint solver we used operates similarly, making it a suitable comparison. In addition, our selection was guided by the intent to utilize representative approaches well-regarded in the field. UMLtoCSP, for instance, is a well-established approach that has been extensively studied in the literature [95]. Similarly, PLEDGE represents a state-of-the-art hybrid constraint-solving approach. This selection ensures good representative approaches for comparison. We executed this experiment on one machine with a 3.2 GHz processor, 32 GB RAM, and a Windows 10 operating system.
@@ 5.2 Discussion on Results
Figure 3 shows the success rates achieved by AVMr, AVMc, AVMrc, AVMo, and RS for all case studies. Table 4 shows the results of the Fisher Exact test and the Wilcoxon test for statistical significance and the results of the effect size calculated using Odds Ratios and Vargha-Delaney’s Â12 measure. For the comparison based on time, Figure 4 shows the time taken for each OCL constraint during 100 evaluations. Table 5 shows the results of the Wilcoxon test for statistical significance and the results of the effect size calculated using Vargha-Delaney’s Â12 measure. The columns in both tables represent a comparison between two algorithms based on statistical significance and effect size corresponding to each case study and its total constraints. The values presented in both tables (i.e., Table 4 and Table 5) refer to the number of times p-value$<$0.05 for the results of Fisher Exact and the Wilcoxon test. Similarly, for the effect size measures (Odds Ratios and Vargha-Delaney’s Â12), the tables show the number of times (including percentage fraction) an algorithm is better than the other one, i.e., Algorithm_1 outperforms ($>$) Algorithm_2. We show aggregated results due to a large number of comparisons among algorithms corresponding to all case studies. Table 6 presents results of comparison among AVMr, AVMc, AVMrc, UMLtoCSP, and PLEDGE based on the percentage of solved MC/DC constraints and the average time.
@@
To compare algorithms based on success rates, the box plots in Figure 3 show that AVMr achieved a median success rate of $\approx$100for five case studies (i.e., GCS, EUR, RnL, SM, and SRS). Similarly, AVMc achieved $\approx$100median success rate for five case studies (i.e., EUR, RnL, SM, SBS, and SRS). AVMo achieved a median success rate of $\approx$100for three case studies (i.e., RnL, SM, and SRS). Whereas RS cannot achieve a 100% median success rate for any of the case studies.
For the statistical comparison between AVMr and AVMc, Table 4 shows that AVMc outperforms AVMr more than 50% of the time for all six case studies. In the case of comparison between AVMr and AVMo, AVMr is better than AVMo for $\approx$60of the time for GCS, EUR, RnL, and SBS case studies. While comparing AVMc and AVMo, AVMc is better than AVMo for more than $\approx$60of the constraints for GCS, EUR, RnL, and SBS case studies. Whereas for two case studies, i.e., SM and SRS, AVMo is better than AVMc and AVMr for $\approx$60of the cases. If we compare AVMr, AVMc, and AVMo with RS, we can notice that all three algorithms outperform RS for more than $\approx$80of the cases for all six case studies.
Comparing AVMr and AVMc based on time, from Figure 4 and Table 5, it is observed that both AVMr and AVMc take approximately an equal amount of time on average. For GCS, RnL, SM, and SRS case studies, AVMc is significantly better than AVMr. For the comparison between AVMr and AVMo, it is observed that AVMo is better for $\approx$60of the time. However, for EUR, SBS, and SRS case studies, AVMr is better than AVMo for $\approx$60of the constraints. In the case of the comparison between AVMc and AVMo, AVMc outperforms AVMo for GCS, SM, and SRS in $\approx$60of the constraints. For the three case studies, i.e., EUR, RnL, and SBS, AVMo outperforms AVMc in $\approx$60of the cases. In the case of comparing AVMr, AVMc, and AVMo with RS, we can analyze that all three algorithms outperform RS for $\approx$50of the cases especially for EUR, RnL, SM, and SBS.
@@
From the box plots shown in Figure 3, it is clear that AVMrc achieved a median success rate of $\approx$100for all six case studies. Whereas individually the AVMr and AVMc achieved $\approx$100median success rate for five case studies. Moreover, AVMo achieved a median success rate of $\approx$100for only three case studies.
@@
For the comparison based on time, the graph in Figure 4 shows that AVMrc takes less time as compared to AVMr, AVMc, and AVMo. The statistical comparison results given in Table 5 show that AVMrc performs better than AVMo for more than $\approx$60of the constraints specifically for GCS, EUR, SBS, and SRS case studies. While comparing AVMrc and AVMc, it is noticed that AVMrc is better than AVMc for more than $\approx$70of the cases specifically for EUR, SBS, and SRS case studies. Whereas for the GCS, RnL, and SM case studies, AVMc is better than AVMrc for $\approx$60of the cases. In the comparison between AVMrc and AVMr, it is observed that AVMrc outperforms AVMr for more than $\approx$90of the time for all six case studies. Finally, for the comparison between AVMrc and RS, it is clear that AVMrc performs better than RS for $\approx$60of the cases specifically for EUR, RnL, SM, and SBS case studies.
@@
After a thorough analysis of UMLtoCSP and PLEDGE results, we observed that the reason for unsolved MC/DC constraints is conflicting clauses. Given that these solvers are not specifically designed for solving MC/DC constraints and handling conflicting constraints, a substantial number of constraints remain unsolved. For instance, PLEDGE operates by taking MC/DC constraints and constructing a syntax tree of these constraints. It then determines which clauses to solve using SMT and which ones with search. During the constraint-solving process, if the clauses to solve using SMT turn out to be conflicting, SMT marks the constraint as unsatisfiable, resulting in unsolved constraints. For example, the MC/DC constraints C1 and C2, as shown in Listing 2, contain conflicting clauses. These conflicts will be identified as unsatisfiable by the SMT and, as a result, will remain unsolved by PLEDGE. Our strategy starts with an MC/DC constraint and assesses it for potential conflicts. If a constraint is conflict-free, our strategy employs CBR or range reduction methods to solve it. In scenarios where a conflict occurs and is deemed unsatisfiable, our strategy bypasses the conflicting constraint and proceeds to the next constraint. This stepwise constraint-solving process systematically addresses individual constraint conflicts and excludes an MC/DC constraint if it is unsolvable.